OpenStudy (anonymous):

If the legs of a right triangle are 8 and 10, find the length of the hypotenuse.

OpenStudy (mathstudent55):

Pythagorean Theorem: \(a^2 + b^2 = c^2\) You have a = 8 and b = 10. Solve for c.

OpenStudy (anonymous):

64+100=164

OpenStudy (mathstudent55):

Ok. That means \(164 = c^2\) or \(c^2 = 164\) Now take the square root of 164 to find c, the length of the hypotenuse.

OpenStudy (anonymous):

80square root of 2

OpenStudy (anonymous):

\[80\sqrt{2}\]

OpenStudy (mathstudent55):

Not quite. To simplify a square root, you need to factor out the largest perfect square. \( \sqrt{164} = \sqrt { 4 \times 41} = 2 \sqrt{41}\)

OpenStudy (anonymous):

fasho good looking thanks for the help

OpenStudy (mathstudent55):

wlcm